[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Jul 19, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Jul 19 20:13:20 UTC 2017


HSU Shift Log for Wed Jul 19 2017 16:13:14 GMT-0400 (EDT)

Shift: H

Forecaster: D Brown

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues: Himawari satellite data is not available in NAWIPS.

Miscellaneous issues: Strong smell of electrical burning developed in  
operations around 1:15 pm. Neal was notified and began to detect the  
source. The smell became stronger and Jann, Patty, Dennis, and the guard  
began enacting emergency plans. The source of the burning could not be  
found and it was decided to sound the fire alarm, evacuate the building,  
and call the fire department. The TWO were ready and sent for 18z. Hugh  
Cobb notified the SDM and Mike contacted CPHC about potential backup. The  
fire department gave the all-clear around 1:45 pm, and everyone was allowed  
back into the building. It did not affect our ability to issue the 21z  
packages. CPHC was notified.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170719/93fd11dd/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list